Name: abort-character Type: integer Default: 0x7 Description:

This is the character that is recognized as the “interrupt” character. The default is control-G. Each character typed is compared to this value, and if it matches, the current action is aborted. While this value is interpreted as a character, it is stored as an integer. Thus to change your abort-character to the character ‘A’, you should set abort-character to 65.

Name: border-width Type: integer Default: 3 Description:

This is the width in pixels of the box that will be drawn around the page image. This box is only drawn if the boolean draw-borders is true.

See Also: draw-borders

Name: cached-pages Type: integer Default: 3 Description:

This is the number of DVI pages that dvitool will cache. A cached page does not have to be reread from the file if it is viewed again, so redisplaying it is fast. However, there is a substantial performance penalty due to swapping if dvitool attempts to cache too many pages, so it behooves you to keep this variable small. You can see the effects of various settings by searching for some non-existant string in a long DVI file while watching various performance indicators (see perfmeter(1)). An empirical test (searching for a non-existant string in a 200 page document) showed an average reduction in total search time of 49 percent when a 2 page cache was used compared with a 10 page cache. A cache size of 2–4 pages seems optimal for most circumstances.

Name: cwd Type: string Description:

This string is the exact analog to the shell’s current working directory. Though it can be altered via the set command, a more appropriate means of changing it is the cd command because cd expects an argument of type filename which can be completed on. To set, cwd is just a string.

See Also: cd, set

Name: draw-borders Type: boolean Default: on Description:

This variable controls whether a border of border-width pixels is drawn around the page image.

Side Effects:

Makes the page image larger, and thus costs some memory (about 4000 bytes for a 3 pixel border). You may wish to turn this off in a memory–starved environment.

See Also: border-width

Name: enable-ansi-keys Type: boolean Default: on Description:

This variable controls the way the function keys are interpreted. When true, the ANSI sequences generated by the function (and left and right) keys are interpreted and functions are run based on bindings that are changeable only at compile time. When false, all the function keys simply print out an innocuous message. N.B. This information is only valid when the function ansi-keys is bound to the string “\e[”.

See Also: ansi-keys

Name: font-path Type: string

Default: /usr/local/fonts/pk

Description:

This is a list of places to look for the font files (more correctly, for pk or pxl files) that dvitool uses. It is in the same format as the PATH of csh(1), i.e. path:path... where each path element may contain the ~ character to reference a user’s home directory or $NAME to reference any environment variable. It is an error to reference an undefined environment variable.

Dvitool looks for both types (PK and PXL) of font files when it goes searching for a font unless the directory of the pathname (one of the path’s above) contains the substring “pk” or the substring “pixel.” If one of those two substrings are found, then dvitool will only look for that type of font file in that directory. So if you have some PXL and some PK files, it makes good sense to use a font-path something like this:

/usr/local/lib/fonts/pk:/usr/local/lib/fonts/pixel

Name: init-cursor-file Type: string Default: null Description:

If set, this string points to a file that contains an image to be used as the default cursor for dvitool. The named file may contain a ~ reference and it is presumed to be the output of iconedit(1). As with all the init- variables, this variable can only be set in the user’s .dvitoolrc file.

See Also: dvitoolrc, init-cursor-xhot, init-cursor-yhot, init-icon-file

Name: init-cursor-xhot Type: integer Default: 9 Description:

This is the horizontal offset in pixels from the upper left corner of the cursor to consider the “hot spot” of the cursor. The hot spot of the cursor is the point considered to be the focal point of the cursor; it is a means of describing which of the 16^{2} pixels in the cursor image is the pixel actually being pointed to. The default is for the circle cursor; it should be changed if you load your own cursor with init-cursor-file.

See Also: init-cursor-file, init-cursor-yhot

Name: init-cursor-yhot Type: integer Default: 9 Description:

The vertical analog to init-cursor-xhot.

See Also: init-cursor-file, init-cursor-xhot Name: init-icon-file Type: string Default: null Description:

This string names a file to be used to create the icon image for dvitool. It is assumed to be the output of iconedit(1).

See Also: init-cursor-file, init-icon-x, init-icon-y

Name: init-icon-x Type: integer Default: 1000 Description:

This is the horizontal position of the pixel at which the upper left hand corner of the dvitool icon will be painted.

See Also: close-window, init-icon-y, init-icon-file

Name: init-icon-y Type: integer Default: 0 Description:

This is the vertical position of the pixel at which the upper left hand corner of the dvitool icon will be painted.

See Also: close-window, init-icon-x, init-icon-file

Name: init-iconic Type: boolean Default: off Description:

When this variable is turned on, dvitool appears in iconic form when it is first painted. This variable provides exactly the same functionality as the suntools(1) flag -Wi. The most common use of init-iconic is when dvitool is invoked from a suntools(1) menu (which doesn’t allow the -Wi flag). When dvitool is started from a .suntools file it is better to specify the -Wi flag on the command line so that when dvitool is invoked some other way, it will come up non-iconic.

See Also: close-window, init-icon-file, init-icon-x, init-icon-y Name: init-scrollbars-on Type: boolean Default: on Description:

This variable controls whether or not dvitool will be created with scrollbars.

Name: init-win-height Type: integer Default: 400 Description:

This is the height in pixels that dvitool will have when it is first created.

See Also: full-screen, init-win-width, zoom-tool

Name: init-win-width Type: integer Default: 1064 Description:

This is the width in pixels that dvitool will have when it is first created. The default value creates a window that is wide enough to see the entire width of an 8.5 X 11 inch page with a 3 pixel wide border and 1 inch margins.

See Also: full-screen, init-win-height, zoom-horizontal

Name: init-win-x Type: integer Default: 88 Description:

This is the horizontal position of the pixel that the upper left hand corner of dvitool will be initially drawn at.

See Also: init-win-y Name: init-win-y Type: integer Default: 200 Description:

This is the vertical position of the pixel that the upper left hand corner of dvitool will be initially drawn at.

See Also: init-win-x

Name: kern-threshold Type: integer Default: 150 Description:

This variable controls how dvitool interprets the stream of DVI commands that it sees as ASCII characters. Recall that a DVI file consists of an arbitrary combination of “set character” commands interspersed with horizontal and vertical movement commands. Dvitool must have some way of deciding whether a horizontal movement is an interword movement, or a horizontal kern. The prior command should be interpreted as an ASCII space, while the latter should not. Kern-threshold is the break–point for that decision: horizontal movements greater than kern-threshold are considered to be ASCII spaces. The value compared with the horizontal movements is actually the width of the widest page mulitplied by the kern-threshold and a constant. Thus for large point sizes or small page widths, kern-threshold may need to be increased. The default value works well for 10 point text.

See Also: ascii-of-selection, extend-selection, line-break-threshold, select-char

Name: left-margin Type: dimension Default: 1in Description:

This is the width of the left margin. Characters may be typeset in the margins, but it is an error to attempt to set a character to the left of the area allocated by left-margin which is “off the page.” Dvitool will automatically make the load image bigger when asked to set any characters in the margins (left, right, top or bottom), but will complain if asked to set a character outside them. If you really want to see those characters, you must change page-height and page-width. As with all variables of type dimension, this value may be specified in either inches (in), scaled points (sp), points (pt), or centimeters (cm).

See Also: page-height, page-width, show-load-image, top-margin Name: line-break-threshold Type: integer Default: 110 Description:

This value is a threshold over which vertical movements will be considered line breaks. A line break is considered a space in the search algorithm. See kernthreshold.

See Also: ascii-of-selection, extend-selection, kern-threshold, select-char

Name: log-filename Type: string Default: null Description:

When this string is non-null, it is the name of a log file in which a copy of all of the messages dvitool produces in it’s message window will be placed. When the file is first openned, it is truncated. It is an error for the user not to be able to open the file. To turn message logging off, set log-file“/ to the null string by entering <return> as the first character.

Name: page-height Type: dimension Default: 11in Description:

The minimum height of the page image. See left-margin.

See Also: left-margin, page-width, top-margin

Name: page-width Type: dimension Default: 8.5in Description:

The minimum width of the page image. See left-margin.

See Also: left-margin, page-height, top-margin Name: show-load-image Type: boolean Default: off Description:

When true, dvitool will draw a 1 pixel wide box around the load image. Each page has its own load image, while there is only 1 global page image. The box reveals exactly which pixels are kept, or cached, for each page.

Name: top-margin Type: dimension Default: 1in Description:

The height of the top margin. See left-margin.

See Also: left-margin, show-load-image